OpenStudy (anonymous):

CHECK MY ANSWER? is the variables in the table a direct variation, an inverse variation, both, or neither? x = 2, 5, 12, 20 y = 30, 12, 5, 3

OpenStudy (anonymous):

i think its a direct variation

OpenStudy (anonymous):

y=2x+2 maybe?

OpenStudy (anonymous):

@dan815 can you help me with this when you get a chance?!

OpenStudy (dan815):

it cannnot be y=2x+2 because it doesnt satisfy the points, what i mean by that is the points they gave you when you sub into your equation, the left side doesnt equal the right side

OpenStudy (skullpatrol):

As x gets larger, y gets smaller, right?

OpenStudy (anonymous):

yes. but wait it is direct variation right??

OpenStudy (skullpatrol):

What is your definition of "direct variation"?

OpenStudy (anonymous):

when they both have a constant change

OpenStudy (dan815):

direct variation is where y=k x means y is propotional to x by a factor of k, if k is positive in this case you can see that when x increases y must also increase

OpenStudy (skullpatrol):

But @veganbee the "direction" of the change matters.

OpenStudy (dan815):

if we think of some equation like this y=10/x now if x is increasing y should be getting smaller because you are dividing 10 by a larger and larger number

OpenStudy (dan815):

^ that is called an inverse variation

OpenStudy (dan815):

since y= k* (1/x)

OpenStudy (anonymous):

ohhhhhh okay i understand that

OpenStudy (dan815):

can you try to figure out your equation

OpenStudy (dan815):

if we believe it is a inverse variation lets try to find our k in y=k/x

OpenStudy (anonymous):

y=60/x .....

OpenStudy (dan815):

ok good

OpenStudy (dan815):

now that makes sense right.. try subbing in the different x values and see if you are getting the corresponding y value

OpenStudy (skullpatrol):



OpenStudy (skullpatrol):